# The Multi-Wavelength Context of Delayed Radio Emission in TDEs: Evidence for Accretion-Driven Outflows

### v1

@@ 2.1 Radio: VLA, MeerKAT, ATCA, and literature data
The full list of literature sources used to compile our radio dataset for each TDE in our sample is given in Appendix B (Table 4). For the purpose of our analysis, we classify each TDE’s radio emission as “Prompt:” first radio detection $<6$ months post-discovery (8 TDEs); “Delayed:” first radio detection $>6$ months post-discovery (12 TDEs), “None:” all radio observations to date resulted in non-detections (5 TDEs), or “Host:” persistent radio emission likely unrelated to the TDE (6 TDEs). The “Host” classification also includes TDEs where the origin of the observed radio emission is ambiguous (e.g., because the emission is too faint to definitively assess variability). We classify TDEs that exhibit multiple radio peaks (ASASSN-15oi and AT2019dsg) based on the onset time of the brightest radio emission episode. The radio classification of each TDE is given in Table 1.
2.2 X-rays: archival data and new Swift-XRT observations
@@
All of the TDEs in our sample have received at least some X-ray follow-up, although the cadence and data quality vary significantly. We provide full details of the literature X-ray observations used in Table 4 in Appendix B. To systematically explore the late-time X-ray properties of our sample, we triggered new target-of-opportunity (TOO) observations with the X-ray Telescope (XRT, Burrows et al. 2005) onboard the Neil Gehrels Swift Observatory (Gehrels et al., 2004) of all TDEs in Table 1 without existing late-time X-ray coverage, which cover the timescale $\sim 2-8$ yrs since discovery. We also re-analyzed all the available Swift-XRT data of TDEs in our sample not already reported in the literature, using the online automated tools released by the Swift-XRT team (Evans et al., 2009) [^2] https://www.swift.ac.uk/user_objects/ and custom scripts (Margutti et al., 2013). We first extracted a 0.3–10 keV count-rate light-curve with a minimum number of counts per bin allowing for dynamic binning. In many cases, these are TDEs with only faint X-rays or non-detections, so we assume the same spectral model for all TDEs for the flux calibration and fix $N_{H}=0$. We considered two spectral models: a constant non-thermal power-law spectrum with photon index $\Gamma=2$; and a constant thermal spectrum consisting of a black-body model with $k_{b}T=0.05\,\rm{keV}$. These spectral models and parameters are broadly representative of the hard (power-law) and soft (black-body) X-ray spectra of TDEs (Saxton et al., 2021; Guolo et al., 2023). For the power-law and black-body models, the average count-to-flux factors are $\approx 5\times 10^{-11}\rm{erg\,s^{-1}cm^{-2}}/\rm{c\,s^{-1}}$ and $\approx 3.3\times 10^{-11}\rm{erg\,s^{-1}cm^{-2}}/\rm{c\,s^{-1}}$, respectively. In the following we employ the power-law model flux calibration for all the TDE X-ray light-curves; we emphasize that the difference of a factor $<2$ between the two flux calibrations has no impact on any of our major conclusions. All newly-analyzed Swift XRT data are presented in Table 6 in Appendix B.
2.3 Optical: Archival spectroscopy and photometry
@@
Optically selected TDEs typically show a common set of spectral features, including hydrogen and/or helium emission lines (Arcavi et al., 2014). All of the TDEs in our sample have optical spectral classifications previously reported in the literature (Foley et al., 2015; Blanchard et al., 2017; Gromadzki et al., 2019; Frederick et al., 2020; van Velzen et al., 2021; Angus et al., 2022; Hammerstein et al., 2023a; Yao et al., 2023). We divide our sample into three spectral types: TDEs with only hydrogen lines in their spectra (hereafter TDE-H), TDEs with both hydrogen and helium lines (TDE-H+He), and TDEs with only helium lines (TDE-He). However, the relative strengths of these lines vary both across TDEs (Arcavi et al., 2014; van Velzen et al., 2021) and as a function of time for specific TDEs (e.g. Nicholl et al. 2019). For this paper, we follow C24 and conservatively define a TDE H+He as a TDE that exhibited He lines at any phase of its evolution, in contrast to some recent literature. By this definition, our sample consists of 5 TDE-H, 24 TDE-H+He, and 2 TDE-He. We list the spectral classifications of our TDEs in Table 1. While most of these are consistent with the classifications reported in van Velzen et al. (2021) and Hammerstein et al. (2023a), three of the TDEs in our sample (ASASSN-14ae, AT2018zr, and AT2018hyz) initially showed only H lines in their spectra, but later developed He II emission at $\sim 50-200$ d post-discovery (Holoien et al., 2014; Short et al., 2020; Hung et al., 2019). We therefore initially classify them as TDE H+He for our analysis, but also consider the possibility that they may more naturally belong to the TDE-H group.
@@
We also collected archival optical and ultraviolet photometry to perform more detailed light curve modeling of the TDEs in our sample, as described in Section 3. We exclude DES14C1kia from this analysis, as the optical dataset for this TDE remains unpublished. A complete listing of the data used and the original references for each dataset are given in Appendix B.
2.4 Host galaxy observations
@@
Most of our TDEs also have well-studied host galaxies. For our analysis in Section 3.1, we use published galaxy stellar masses ($M_{\rm gal}$), generally calculated via broadband modeling of archival near-IR through UV photometry, and published stellar velocity dispersions ($\sigma*$) derived from high-resolution optical spectroscopy (Wevers et al., 2017, 2019a; Yao et al., 2023). We also use the optical $u-r$ colors of the TDE host galaxies in our sample as compiled by French et al. (2020) and references therein (van Velzen et al., 2021; Hammerstein et al., 2021; Yao et al., 2023).
3 Multi-Wavelength Analysis
@@
3.1 Optical spectroscopic classification, host properties, and SMBH masses
@@
We first search for possible correlations between our radio classifications and the basic optical properties of the TDEs in our sample and their host galaxies. We find that 4 of the 12 TDEs that exhibit delayed radio emission have an optical spectral classification of TDE-H. For our sample of 31 objects containing 5 TDE-H total, the probability of at least 4 of the 5 TDEs also falling into this group by chance is $p=0.06$. Thus, based on the current dataset and our conservative definition of a TDE-H as one without any He emission at any epoch, we find no statistically significant correlation between the timing of radio emission and optical spectral type. However, interestingly, if the three TDEs for which He lines only emerge at $t>50$ d are reclassified as TDE-H rather than TDE-H+He, then our sample has a total of 8 TDE-H, of which 7 exhibit delayed radio emission. In this case, the probability of chance coincidence decreases to $p=0.002$.
@@
Previous studies of the optical $u-r$ colors of TDE host galaxies have shown that TDE hosts are preferentially located in the “green valley” (French et al. 2020 and references therein; Hammerstein et al. 2021; Yao et al. 2023). We find no obvious correlation between the bulk host galaxy classifications (quiescent vs. star forming; red vs. green vs. blue) and the type of radio emission in our TDE sample (prompt vs. delayed vs. none/host). We defer further investigation of possible impacts of the host environment on the radio properties of TDEs to future work.
@@
Finally, we compute the SMBH masses for all TDEs in our sample. As SMBH masses previously reported in the literature are derived using several different methods, here we use two independent techniques: a transient-independent method based on observations of the TDE host galaxies ($M_{\rm BH,host}$) and a TDE-only method based on fitting the TDE optical and UV light curves ($M_{\rm BH,mosfit}$). First, we use archival host galaxy observations to self-consistently recalculate the SMBH masses for all of our TDEs and present them in Table 2, along with the original measurement used to derive each value. For galaxies with published stellar velocity dispersions, we calculate the black hole mass using the $M_{\rm BH}-\sigma*$ relation from Kormendy & Ho (2013). For galaxies that do not have published velocity dispersions but do have published stellar masses, we use the $M_{\rm BH}-M_{\rm gal}$ relation for TDE host galaxies derived by Yao et al. (2023), which is calibrated against the Kormendy & Ho (2013) relation. [^3] OGLE17aaj’s host galaxy does not have a published $\sigma*$ measurement or $M_{\rm gal}$ value; we therefore use the SMBH mass estimate from Gromadzki et al. (2019), $\log(M_{\rm BH}/M_{\odot})=7.37$.
@@
TDE light curves have also been used to estimate SMBH masses directly (Mockler et al., 2019; Ryu et al., 2020; Sarin & Metzger, 2024; Mummery et al., 2024b). The masses derived in this way are typically found to be consistent with SMBH masses derived from the host galaxy properties, with some scatter. We choose to model the optical and UV light curves of our TDEs using the Modular Open Source Fitter for Transients (MOSFiT; Guillochon et al. 2018; Mockler et al. 2019) to obtain a second, independent estimate of the SMBH mass for each TDE. For our sample, we find that the two methods yield generally consistent results, although on average the MOSFiT-derived masses are slightly higher (Figure 1, left). This is similar to the results obtained by Mockler et al. (2019) for their TDE sample. For our analysis, we do not make any claims as to which method of deriving the SMBH mass is more accurate; instead, we consider both sets of masses independently whenever possible. We discuss our MOSFiT modeling further in the next two sections.
@@
Table 2: SMBH masses for the TDEs in our sample derived from properties of their host galaxies. The $M_{\rm BH,host}$ values have been computed using the $M_{\rm BH}-\sigma*$ relation of Kormendy & Ho (2013) when possible, or using the $M_{\rm gal}-M_{\rm BH}$ relation from Yao et al. (2023) otherwise. The source used for each galaxy stellar mass or $\sigma*$ measurement is indicated in the Reference column.
@@
ASASSN-14ae $53\pm 2$ Wevers et al. (2019b) $6.0\pm 0.3$
DES14C1kia $10.1$ Wevers et al. (2019b) $6.6$
ASASSN-14li $81\pm 2$ Wevers et al. (2019b) $6.8\pm 0.3$
iPTF15af $106\pm 2$ Wevers et al. (2019b) $7.3\pm 0.3$
ASASSN-15oi $61\pm 7$ Wevers et al. (2019b) $6.2\pm 0.4$
iPTF16axa $82\pm 3$ Wevers et al. (2019b) $6.8\pm 0.3$
@@
AT2018hco $10.03^{+0.12}_{-0.16}$ Hammerstein et al. (2023a) $6.5\pm 0.3$
AT2018hyz $66.62\pm 3.12$ Hammerstein et al. (2023b) $6.4\pm 0.3$
AT2018lna $36.43\pm 4.52$ Hammerstein et al. (2023b) $5.2\pm 0.4$
ASASSN-19bt $10.04^{+0.34}_{-0.04}$ Holoien et al. (2019a) $6.5\pm 0.4$
AT2019azh $68.01\pm 2.02$ Hammerstein et al. (2023b) $6.4\pm 0.3$
@@
3.2 MOSFiT modeling: physical parameters
@@
We fit the optical and UV photometry for the TDEs in our sample with MOSFiT (Guillochon et al., 2018), using version 1.1.9 of the TDE module first developed by Mockler et al. (2019) following Guillochon & Ramirez-Ruiz (2013); Guillochon et al. (2014). The original references for the included data are given in Table 4 in Appendix B. We use the same setup, priors, and model assumptions as those described in Nicholl et al. (2022). In particular, MOSFiT assumes that mass falling onto the SMBH accretes promptly, such that the observed optical luminosity is directly proportional to the mass fallback rate and the accretion rate, $L=\epsilon\dot{M}c^{2}$ (where $\epsilon$ is the radiative efficiency). [^4] A possible viscous delay is included in the model, but when applied to our observations this delay is generally found to be very short (consistent with zero in most cases), in line with previous work (Mockler et al., 2019; Mockler & Ramirez-Ruiz, 2021; Nicholl et al., 2022). This corresponds to Case 1 as defined in the introduction. This assumption likely breaks down at very late times, when TDEs’ optical/UV emission often exhibits a plateau phase that is not well-captured by MOSFiT but can be better modeled by a bare accretion disk (van Velzen et al., 2019a; Mummery et al., 2024a, b). For the TDEs in our sample that show optical/UV plateaus, we therefore only include photometry prior to the onset of the plateau phase in our modeling, as detailed in Appendix B.
We are able to find a converged model solution for all TDEs in our sample except for DES141Ckia (which has no published photometry) and AT2018fyk (whose double-peaked light curve is not well-modeled by MOSFiT’s one-zone model assumptions). A complete description of the free parameters of our model, together with the best-fit values and $1\sigma$ confidence intervals for the physically relevant parameters are given in Appendix C. We first use these results to search for possible correlations between the radio properties of our TDEs and other physical parameters of the disruption. We divide our sample into three groups for this analysis: those with radio type “Prompt” in Table 1 (i.e., TDEs that show early/prompt radio emission), “Delayed” TDEs in Table 1 (those that show late/delayed radio emission), and “None” or “Host” TDEs in Table 1 (those that show no radio emission or only host or faint ambiguous radio emission). We also considered the None and Host TDEs as two independent categories, but this does not significantly change our results. We then performed an Anderson-Darling (AD) test [^5] The Anderson-Darling test is similar to the Kolmogorov–Smirnov (KS) test used by some previous authors (e.g., Nicholl et al. 2022), but it has been shown to have added sensitivity to the tails of the distributions in datasets with large numbers of free parameters (Hou et al., 2009). We note that a KS test applied to our data gives similar results. for each of the parameters listed in Appendix C (Table 7) plus $M_{\rm BH,host}$ (the SMBH masses derived from host galaxy properties presented in Table 2).

### v2

@@ II.1 Radio: VLA, MeerKAT, ATCA, and literature data
The full list of literature sources used to compile our radio dataset for each TDE in our sample is given in Appendix B (Table 6). For the purpose of our analysis, we classify each TDE’s radio emission as “Prompt:” first radio detection <6<6 months post-discovery (8 TDEs); “Delayed:” first radio detection >6>6 months post-discovery (12 TDEs), “None:” all radio observations to date resulted in non-detections (5 TDEs), or “Host:” persistent radio emission likely unrelated to the TDE (6 TDEs). The “Host” classification also includes TDEs where the origin of the observed radio emission is ambiguous (e.g., because the emission is too faint to definitively assess variability). We classify TDEs that exhibit multiple radio peaks (ASASSN-15oi and AT2019dsg) based on the onset time of the brightest radio emission episode. We note that some TDEs in our sample are significantly less well-observed in the radio than others. In particular, several of the TDEs that we classify as radio type Delayed have no radio observations at $t<6$ months, meaning that we have no way of knowing if they also displayed prompt radio emission and are thus misclassified here, potentially diluting any underlying trends (Section IV.2.3). The radio classification of each TDE is given in Table 1.
II.2 X-rays: archival data and new Swift-XRT observations
@@
All of the TDEs in our sample have received at least some X-ray follow-up, although the cadence and data quality vary significantly. We provide full details of the literature X-ray observations used in Table 6 in Appendix B. To systematically explore the late-time X-ray properties of our sample, we triggered new target-of-opportunity (TOO) observations with the X-ray Telescope (XRT, Burrows et al. 2005) onboard the Neil Gehrels Swift Observatory (Gehrels et al., 2004) of all TDEs in Table 1 without existing late-time X-ray coverage, which cover the timescale $\sim 2-8$ yrs since discovery. We also re-analyzed all the available Swift-XRT data of TDEs in our sample not already reported in the literature, using the online automated tools released by the Swift-XRT team (Evans et al., 2009) [^2] https://www.swift.ac.uk/user_objects/ and custom scripts (Margutti et al., 2013). We first extracted a 0.3–10 keV count-rate light-curve with a minimum number of counts per bin allowing for dynamic binning. In many cases, these are TDEs with only faint X-rays or non-detections, so we assume the same spectral model for all TDEs for the flux calibration and fix $N_{\rm H,int}=0$ (meaning that we assume zero $N_{\rm H}$ in the TDE host galaxy and we account only for the Galactic $N_{\rm H}$ expected along the line of sight). We considered two spectral models: a constant non-thermal power-law spectrum with photon index $\Gamma=2$; and a constant thermal spectrum consisting of a black-body model with $k_{b}T=0.05\,\rm{keV}$. These spectral models and parameters are broadly representative of the hard (power-law) and soft (black-body) X-ray spectra of TDEs (Saxton et al., 2021; Guolo et al., 2024). For the power-law and black-body models, the average count-to-flux factors are $\approx 5\times 10^{-11}\rm{erg\,s^{-1}cm^{-2}}/\rm{c\,s^{-1}}$ and $\approx 3.3\times 10^{-11}\rm{erg\,s^{-1}cm^{-2}}/\rm{c\,s^{-1}}$, respectively. In the following we employ the power-law model flux calibration for all the TDE X-ray light-curves; we emphasize that the difference of a factor <2<2 between the two flux calibrations has no impact on any of our major conclusions. All newly-analyzed Swift XRT data are presented in Table 8 in Appendix B.
II.3 Optical: Archival spectroscopy and photometry
Optically selected TDEs typically show a common set of spectral features, including hydrogen and/or helium emission lines (Arcavi et al., 2014). The relative strengths of these emission lines reveal the ionization state of the stellar debris cloud (Roth et al., 2016) and may provide clues to the underlying physics of the disruption; for example, Nicholl et al. (2022) found that TDEs with only hydrogen in their spectra may be either incomplete stellar disruptions or cases where accretion disk formation is delayed.
@@ II.3 Optical: Archival spectroscopy and photometry
All of the TDEs in our sample have optical spectral classifications previously reported in the literature (Foley et al., 2015; Blanchard et al., 2017; Gromadzki et al., 2019; Frederick et al., 2020; van Velzen et al., 2021; Angus et al., 2022; Hammerstein et al., 2023b; Yao et al., 2023). We divide our sample into three spectral types: TDEs with only hydrogen lines in their spectra (hereafter TDE-H), TDEs with both hydrogen and helium lines (TDE-H+He), and TDEs with only helium lines (TDE-He). However, the relative strengths of these lines vary both across TDEs (Arcavi et al., 2014; van Velzen et al., 2021) and as a function of time for specific TDEs (e.g., Nicholl et al. 2019). For this paper, we follow common practice (e.g., van Velzen et al. 2021; Hammerstein et al. 2023b) and define each TDE’s spectral class based on the lines visible near peak optical light. By this definition, our sample consists of 8 TDE-H, 21 TDE-H+He, and 2 TDE-He. We list the spectral classifications of our TDEs in Table 1. While most of these are consistent with the classifications reported in C24, who conservatively define a TDE H+He as a TDE that exhibited He lines at any phase of its evolution, three of the TDEs in our sample (ASASSN-14ae, AT2018zr, and AT2018hyz) initially showed only H lines in their spectra, but later developed He II emission at $\sim 50-200$ d post-discovery (Holoien et al., 2014; Short et al., 2020; Hung et al., 2019). While C24 classify these three transients as TDE-H+He, we choose to retain a TDE-H classification for them. This is the best choice to ensure consistency across our sample because 4 of our remaining 5 TDE-H have no published optical spectra at $t>50$ d, so it remains possible that they could have developed He lines at late times as well.
@@
We also collected archival optical and ultraviolet photometry to perform more detailed light curve modeling of the TDEs in our sample, as described in Section III. We exclude DES14C1kia from this analysis, as the optical dataset for this TDE remains unpublished. A complete listing of the data used and the original references for each dataset are given in Appendix B.
II.4 Host galaxy observations
@@
Most of our TDEs also have well-studied host galaxies. For our analysis in Section III.1, we use published galaxy stellar masses ($M_{\rm gal}$) calculated via broadband modeling of archival near-IR through UV photometry and published stellar velocity dispersions ($\sigma*$) derived from high-resolution optical spectroscopy (Wevers et al., 2017, 2019b; Yao et al., 2023). We also use the optical uru-r colors of the TDE host galaxies in our sample as compiled by French et al. (2020) and references therein (van Velzen et al., 2021; Hammerstein et al., 2021; Yao et al., 2023).
III Multi-Wavelength Analysis
@@
III.1 Optical spectroscopic classification, host properties, and SMBH masses
@@
We first search for possible correlations between our radio classifications and the basic optical properties of the TDEs in our sample and their host galaxies. We find that 7 of the 12 TDEs that exhibit delayed radio emission have an optical spectral classification of TDE-H. For our sample of 31 objects containing 8 TDE-H total, the probability of at least 7 of the 8 TDEs also falling into this group by chance is $p=0.002$. This suggests a correlation between a TDE’s optical spectral type and the timing of its radio emission at the $\sim 3\sigma$ level. We discuss the implications of this further in Section IV.
@@
Previous studies of the optical uru-r colors of TDE host galaxies have shown that TDE hosts are preferentially located in the “green valley” (French et al. 2020 and references therein; Hammerstein et al. 2021; Yao et al. 2023). We find no obvious correlation between the bulk host galaxy classifications (quiescent vs. star forming; red vs. green vs. blue) and the type of radio emission in our TDE sample (prompt vs. delayed vs. none/host). We defer further investigation of possible impacts of the host environment on the radio properties of TDEs to future work.
@@
Finally, we compute the SMBH masses for all TDEs in our sample. As SMBH masses previously reported in the literature are derived using several different methods, here we use two independent techniques: a transient-independent method based on observations of the TDE host galaxies ($M_{\rm BH,host}$) and a TDE-only method based on fitting the TDE optical and UV light curves ($M_{\rm BH,mosfit}$). First, we use archival host galaxy observations to self-consistently recalculate the SMBH masses for all of our TDEs and present them in Table 2, along with the original measurement used to derive each value. For galaxies with published stellar velocity dispersions, we calculate the black hole mass using the $M_{\rm BH}-\sigma*$ relation from Kormendy and Ho (2013). For galaxies that do not have published velocity dispersions but do have published stellar masses, we use the $M_{\rm BH}-M_{\rm gal}$ relation for TDE host galaxies derived by Yao et al. (2023), which is calibrated against the Kormendy and Ho (2013) relation. [^3] OGLE17aaj’s host galaxy does not have a published $\sigma*$ measurement or $M_{\rm gal}$ value; we therefore use the SMBH mass estimate from Gromadzki et al. (2019), $\log(M_{\rm BH}/M_{\odot})=7.37$.
@@
TDE light curves have also been used to estimate SMBH masses directly (Mockler et al., 2019; Ryu et al., 2020b; Sarin and Metzger, 2024; Mummery et al., 2024). The masses derived in this way are typically found to be consistent with SMBH masses derived from the host galaxy properties, with some scatter. [^4] While a detailed intercomparison of the different light curve models is beyond the scope of this work, we refer the interested reader to Table 7 and the discussion in Section 7 of Hammerstein et al. (2023b) for a comparison of the SMBH masses computed using two of these light curve-based models, MOSFiT (Mockler et al., 2019) and TDEMass (Ryu et al., 2020b). We choose to model the optical and UV light curves of our TDEs to obtain a second, independent estimate of the SMBH mass for each TDE using the most widely-used of these three light curve-based models: the Modular Open Source Fitter for Transients (MOSFiT; Guillochon et al. 2018; Mockler et al. 2019). For our sample, we find that the host and MOSFiT methods yield generally consistent results, although on average the MOSFiT-derived masses are slightly higher (Figure 1, left). This is similar to the results obtained by Mockler et al. (2019) for their TDE sample. For our analysis, we do not make any claims as to which method of deriving the SMBH mass is more accurate; instead, we consider both sets of masses independently whenever possible. We discuss our MOSFiT modeling further in the next two sections.
@@
Table 2: SMBH masses for the TDEs in our sample derived from properties of their host galaxies. The $M_{\rm BH,host}$ values have been computed using the $M_{\rm BH}-\sigma*$ relation of Kormendy and Ho (2013) when possible, or using the $M_{\rm gal}-M_{\rm BH}$ relation from Yao et al. (2023) otherwise. The source used for each galaxy stellar mass or $\sigma*$ measurement is indicated in the Reference column.
@@
ASASSN-14ae $53\pm 2$ Wevers et al. (2019a) $6.0\pm 0.3$
DES14C1kia $10.1$ Wevers et al. (2019a) $6.6$
ASASSN-14li $81\pm 2$ Wevers et al. (2019a) $6.8\pm 0.3$
iPTF15af $106\pm 2$ Wevers et al. (2019a) $7.3\pm 0.3$
ASASSN-15oi $61\pm 7$ Wevers et al. (2019a) $6.2\pm 0.4$
iPTF16axa $82\pm 3$ Wevers et al. (2019a) $6.8\pm 0.3$
@@
AT2018hco $10.03^{+0.12}_{-0.16}$ Hammerstein et al. (2023b) $6.5\pm 0.3$
AT2018hyz $66.62\pm 3.12$ Hammerstein et al. (2023a) $6.4\pm 0.3$
AT2018lna $36.43\pm 4.52$ Hammerstein et al. (2023a) $5.2\pm 0.4$
ASASSN-19bt $10.04^{+0.34}_{-0.04}$ Holoien et al. (2019b) $6.5\pm 0.4$
AT2019azh $68.01\pm 2.02$ Hammerstein et al. (2023a) $6.4\pm 0.3$
@@
III.2 MOSFiT modeling: physical parameters
@@
We fit the optical and UV photometry for the TDEs in our sample with MOSFiT (Guillochon et al., 2018), using version 1.1.9 of the TDE module first developed by Mockler et al. (2019) following Guillochon and Ramirez-Ruiz (2013); Guillochon et al. (2014). MOSFiT assumes that mass falling onto the SMBH accretes promptly, such that the observed optical luminosity is directly proportional to both the mass fallback rate and the accretion rate and we can define $L=\epsilon\dot{M}c^{2}$ (where $\epsilon$ is the radiative efficiency). [^5] A possible viscous delay is included in the model, but when applied to our observations this delay is generally found to be very short (consistent with zero in most cases), in line with previous work (Mockler et al., 2019; Mockler and Ramirez-Ruiz, 2021; Nicholl et al., 2022). This corresponds to Case 1 as defined in the introduction. This assumption likely breaks down at very late times, when TDEs’ optical/UV emission often exhibits a plateau phase that is not well-captured by MOSFiT but can be better modeled by a bare accretion disk (van Velzen et al., 2019b; Mummery et al., 2025, 2024). For the TDEs in our sample that show optical/UV plateaus, we therefore only include photometry prior to the onset of the plateau phase in our modeling. The original references for the included data, together with the photometry cutoff times where applicable, are given in Table 6 in Appendix B.
The MOSFiT TDE model has nine free physical parameters (Mockler et al., 2019; Nicholl et al., 2022), of which eight are of potential interest for discriminating among competing TDE emission models: the viscous delay time $t_{v}$, the disruption impact parameter $b$, the SMBH mass $M_{\rm BH,mosfit}$, the column density along the line of sight in the host galaxy $n_{\rm H,host}$, the normalization of the photospheric radius $R_{\rm ph,0}$ at the time when the observed bolometric luminosity of the TDE reaches $L_{\rm Edd}$, the corresponding power law exponent $l_{\rm ph}$ that relates the photospheric radius to the instantaneous luminosity as $R_{\rm ph}\propto R_{\rm ph,0}(L/L_{\rm Edd})^{l_{\rm ph}}$, the mass of the disrupted star $M_{*}$, and the efficiency with which accreting mass is converted to radiation, $\epsilon$. The remaining parameter is the time of first fallback relative to the first data point in the light curve, $t_{0}$. Although not directly physically relevant, $t_{0}$ can be used in combination with the light curves to calculate the date of disruption, MJD0, and the number of rest-frame days between disruption and peak optical/UV light, $t_{\rm rise}$ (which has been suggested to correlate with other physical properties of TDEs; e.g., Rees 1988; van Velzen et al. 2021). We derive two additional parameters of possible physical interest from our MOSFiT solutions: $\dot{M}_{\rm max}$, the maximum accretion rate onto the SMBH during each TDE, and $L_{\rm max}$, the maximum bolometric luminosity during the TDE. For our analysis, we normalize $\dot{M}_{\rm max}$ by $\dot{M}_{\rm Edd}$ (as defined in Section III.3) to remove the dependence on the efficiency.
Using the same setup, priors, and model assumptions as those described in Nicholl et al. (2022), we are able to find a converged model solution for all TDEs in our sample except for DES141Ckia (which has no published photometry) and AT2018fyk (whose double-peaked light curve is not well-modeled by MOSFiT’s one-zone model assumptions). The best-fit physically relevant parameters and $1\sigma$ confidence intervals for our fits are given in Table 3. We note that our results for the TDEs also studied by Nicholl et al. (2022) are similar to those reported in that work.